OpenStudy (anonymous):

Find the inverse of the function. f(x) = 4x3 - 7 f(x) = 7x + 2 f(x) = x3 - 3

OpenStudy (cwrw238):

put y = f(x) then find x in terms of y replace y by x to give inverse function f-1(x)

OpenStudy (cwrw238):

eg y = 4x^3 - 7 4x^3 = y - 7 x^3 = (y-7) / 4 now find x

OpenStudy (anonymous):

thank you